Circular No. 5744

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER-LEVY (1993e)

Further selected precise positions:

1993 UT α_{2000} δ_{2000} m_{1} Observer
Mar. 15.57153 12^{h}30^{m}52.^{s}17 −4^{o}28′12′′.3 16.0 Endate
15.58611 12 30 51.73 −4 28 09.7  _{”}
26.23333 12 25 43.71 −3 58 04.9 15  Naranjo
28.24097 12 24 44.97 −3 52 16.3  _{”}
28.25677 12 24 44.36 −3 52 13.4  Tatum
29.53692 12 24 06.74 −3 48 25.0  Urata
29.70046 12 24 01.71 −3 47 56.4  _{”}
Apr. 1.88329 12 22 28.62 −3 38 37.4 13.1 Meyer
1.88656 12 22 28.71 −3 38 36.6 13.8  _{”}

K. Endate (Kitami). 0.25-m f/2.6 Schmidt camera. Measurer K. Watan-

abe. Prediscovery observations. O. Naranjo (Mérida). 1.0-m Schmidt. Independent discovery. J. B. Tatum and D. D. Balam (Victoria). 0.25-m Schmidt. T. Urata (Oohira). 0.25-m f/3.4 hyperboloid astrocamera + CCD. E. Meyer and H. Raab (Linz). 0.3-m f/5.2 Schmidt-Cassegrain + CCD.

Attempts at orbit determination from the Mar. 15-Apr. 1 arc suggest that a parabolic solution is no longer viable, but the likely location of the object near the surface of Jupiter’s sphere of influence—not to mention the near impossibility of measuring the object’s center of mass—continues to make the unequivocal determination of an elliptical orbit extremely difficult. A very close encounter with Jupiter during 1992 continues to be a distinct possibility, the most viable such solutions having still smaller orbital eccentricities than in the example on IAUC 5726 and the Jupiter encounter earlier in the year. The heliocentric orbit below puts the object 0.007 AU from Jupiter on 1992 May 16 (tidal breakup presumably requiring an approach to 0.001 AU) and indicates that the object is at least temporarily in orbit about Jupiter and currently near apojove at a distance of 0.31 AU.

Epoch = 1993 Apr. 3.0 TT

T = 1997 Sept. 4.494 TT  ω = 348.^{o}427
e = 0.07169  Ω = 343.394 }2000.0
q = 4.71659 AU  i = 2.206
a = 5.08085 AU n^{o} = 0.086060 P = 11.45 years

1993 April 3 Brian G. Marsden